Presented here are the comparative balance sheets of Hames, Inc., at December 31, 2017 and 2016. Sales for the year ended December 31, 2017, totaled $640,000. HAMES, INC., Balance Sheets December 31, 2017 and 2016 2017 2016 Assets Cash $ 20,000 $ 20,000 Accounts receivable 78,000 72,000 Merchandise inventory 103,000 99,000 Total current assets $ 201,000 $ 191,000 Land 50,000 40,000 Plant and equipment 125,000 110,000 Less: Accumulated depreciation (65,000 ) (60,000 ) Total assets $ 311,000 $ 281,000 Liabilities Short-term debt $ 18,000 $ 17,000 Accounts payable 64,400 75,500 Other accrued liabilities 20,000 18,000 Total current liabilities $ 102,400 $ 110,500 Long-term debt 22,000 30,000 Total liabilities $ 124,400 $ 140,500 Stockholders’ Equity Common stock, no par, 100,000 shares authorized 40,000 and 25,000 shares issued, respectively $ 74,000 $ 59,000 Retained earnings: Beginning balance $ 81,500 $ 85,000 Net income for the year 51,100 1,500 Dividends for the year (20,000 ) (5,000 ) Ending balance $ 112,600 $ 81,500 Total stockholders’ equity $ 186,600 $ 140,500 Total liabilities and stockholders’ equity $ 311,000 $ 281,000 Required: a. Calculate ROI for 2017. (Do not round intermediate calculations. Round your final answer to 2 decimal places.) b. Calculate ROE for 2017. (Round your answer to 1 decimal place.) c. Calculate working capital at December 31, 2017. d. Calculate the current ratio at December 31, 2017. (Round your answer to 2 decimal places.) e. Calculate the acid-test ratio at December 31, 2017. (Round your answer to 2 decimal places.) f. Assume that on December 31, 2017, the treasurer of Hames, Inc., decided to pay $15,000 of accounts payable. What impact, if any, this receipt will have on the answers you calculated for requirements a-d (increase, decrease, or no effect) a. ROI for the year ended December 31, 2017: Increase No effect Decrease b. ROE for the year ended December 31, 2017: Increase Decrease No effect c.Working capital as at December 31, 2017: Increase No effect Decrease d. Current ratio as at December 31, 2017: Increase No effect Decrease g. Assume that instead of paying $15,000 of accounts payable on December 31, 2017. Hames, Inc., collected $15,000 of accounts receivable. What impact, if any, this receipt will have on the answers you calculated for requirements a-d (increase, decrease, or no effect) a. ROI for the year ended December 31, 2017: Decrease No effect Increase b. ROE for the year ended December 31, 2017: Increase Decrease No effect c. Working capital as at December 31, 2017: No effect Increase Decrease d. Current ratio as at December 31, 2017: Increase No effect Decrease
Answer #1

1) ROI = Margin × Turnover

= (Net income / Sales) × (Sales / Average total assets)

= (51,100 / 640,000) X (640,000 / [(311,000 + 281,000) / 2]

= 17.26%

 

2) ROE = Net income / Average stockholders' equity

= 51,100 / [(186,600 $ 140,500) / 2]

= 31.24%

3) Working capital = Current assets – Current liabilities

= $201,000 - $102,400

= 98,600

4) Current ratio = Current assets / Current liabilities = 201,000 / 102400 = 1.96

5) Acid test ratio = (Cash + Accounts receivable) / Current liabilities

= (20,000 + 78,000) / 102,400

= 0.96

Have answered more than 4 parts